# Ajad Chhatkuli1, Xuanjing Huang5, Yu-Gang Jiang5, Luc Van Gool1, Danda Pani Paudel1 1

### v1

@@ 3.2 EgoNight-VQA Benchmark Reconstruction
QA Task Taxonomy. To thoroughly assess models from multiple perspectives, we define a diverse taxonomy of 12 QA tasks. Some of these categories are well-studied and have been explored in previous egocentric VQA benchmarks, such as object/action/text recognition, counting, and spatial reasoning. Others are much less studied or newly proposed in EgoNight-VQA, including scene sequence and navigation (which require not only visual perception but also memory and spatial reasoning), illumination recognition and illumination change (designed to test models’ understanding of lighting concepts), and non–common-sense reasoning (e.g., detecting abnormal cases such as a door inserted into a wall in the synthetic data). More detailed explanations of QA types can be found in the Appendix Sec. A.4.1. We further organize these categories into paired and unpaired QA types, depending on whether the same questions can be consistently applied across day–night counterparts:
  • Paired QA Types. These cover contexts that remain unchanged across day and night, allowing the same QA pairs to be used for both videos and thus providing a clean testbed for measuring performance gaps. Specifically, we include: ① object recognition, ② text recognition, ③ spatial reasoning, ④ scene sequence, ⑤ navigation, ⑥ counting of static, ⑦ action recognition, and ⑧ non–common-sense reasoning.
  • Unpaired QA Types. These include categories that are impractical to pair across day and night, or are only meaningful in the nighttime condition. We consider: ① lighting recognition, ② lighting dynamic, ③ dynamic detection, and ④ counting of dynamic. We control QA clip duration by task type. For static or spatial tasks (e.g., object recognition, lighting recognition), we use short clips of 3 seconds to minimize redundancy; For dynamic or temporal tasks (e.g., action recognition, navigation), the entire video is used to capture the complete context. Following recent works [13, 6], we adopt the open-ended QA setting over the closed-form multiple-choice format, as it better reflects natural human–AI interactions.
@@
All three stages are powered by GPT-4.1. Empirically, we find that both the QA-type-specific prompting and the inclusion of daytime videos substantially improve the quality and reliability of the generated QA pairs.
@@
Dataset Statistics. EgoNight-VQA comprises 3,658 high-quality, fully human-verified QA pairs across 12 task types, sourced from EgoNight-Synthetic, EgoNight-Sofia, and EgoNight-Oxford, with an average of 40 pairs per video. Detailed statistics on QA distribution, video durations, task difficulties, scenarios, and illumination are shown in Fig. 4. Overall, EgoNight-VQA provides a diverse and comprehensive benchmark for evaluating egocentric vision models under the challenging nighttime conditions.
@@ 3.3 Benchmarks Beyond Egocentric VQA
Day-Night Correspondence Retrieval. To further assess model capabilities beyond VQA, we introduce day–night correspondence retrieval, which evaluates a model’s ability to match visual content across illumination conditions. Specifically, we define two subtasks: i) Spatial Retrieval (Scene Recognition). Spatial retrieval, or scene recognition, is a long-standing vision task [75, 76, 77]. Here, it is extended: given a query clip and a set of $N$ candidate clips of equal duration $s$, the model must retrieve the one depicting the same scene. This evaluates a model’s ability to capture and relate spatial relations in egocentric videos, e.g., distinguishing a bedroom from a bathroom or another bedroom. We built this benchmark with 1000 randomly generated meta-tasks. Each task samples a query clip, and the candidate set includes its temporally aligned counterpart (with a temporal shift for added difficulty) plus N1N-1 negatives from other scenes. Performance is measured by Top-1 accuracy across all tasks. In our setup, we use $N=10$, $s=10$ seconds, and a temporal shift of [10, 20] frames. Both Day (query) $\rightarrow$ Day (database) and Day $\rightarrow$ Night settings are evaluated.
@@
ii) Temporal Localization. We further design a temporal localization task to test whether models can align video segments across dynamics. Given a query clip of duration $s$, the model must localize it within the corresponding full video by predicting its start and end timestamps $(t_{i},t_{j})$, directly evaluating temporal reasoning (e.g., grounding “The door is being locked” to 10–20s). We construct 1000 meta-tasks, each generated by randomly sampling one clip from its parent full video that is also randomly selected. Inspired by temporal grounding literature [78], we adopt mean Intersection-over-Union (mIoU) between the predicted interval $(t_{i},t_{j})$ and the ground-truth interval $(t_{i}^{*},t_{j}^{*})$ as the evaluation metric. Consistent with spatial retrieval, we set $s=10$ seconds and evaluate both Day $\rightarrow$ Day and Night $\rightarrow$ Day settings.
@@
Egocentric Depth Estimation at Night. Depth estimation is a fundamental component of computer vision. On the one hand, extensive research [79, 80, 81] has focused on depth estimation in non-egocentric settings (typically not with fisheye cameras), while egocentric depth estimation remains largely underexplored, especially under nighttime conditions. On the other hand, recent works [82, 83] suggest that incorporating depth can enhance models’ spatial reasoning abilities. These two observations motivate us to construct an auxiliary benchmark for egocentric depth estimation at night. Specifically, we use EgoNight-Synthetic as the testbed, where ground-truth depth maps are provided by the rendering engine. Thanks to the day–night aligned design, we can quantitatively evaluate models under both controlled daytime and nighttime conditions. For evaluation, we adopt standard depth estimation metrics, including absolute relative error (AbsRel), $\delta_{1}(1.25)$, $\delta_{2}(1.25^{2})$, and $\delta_{3}(1.25^{3})$, where $\delta_{k}$ measures the percentage of predicted pixels whose relative error is within a threshold of $1.25^{k}$.
@@ 4.1 Evaluated MLLMs & Metrics
We evaluate a broad set of state-of-the-art MLLMs on the proposed benchmarks. i) For EgoNight-VQA, we include two closed-source commercial models, GPT-4.1 [36] and Gemini 2.5 Pro [37]; eight open-source models, Qwen2.5-VL (3B, 7B, 72B) [31], VideoLLaMA3 (7B) [33], InternVL3 (8B) [32], GLM-4.1V (9B-Base) [35], and LLaVA-NeXT-Video (7B) [34]; as well as EgoGPT [1], one of the few open-source egocentric models tailored for open-ended QA. Following prior work [13, 25], we adopt an LLM-as-a-Judge strategy to assess semantic consistency between predictions and ground truth, and report average accuracy across the test sets. ii) For day–night correspondence retrieval, we benchmark feature-based retrieval methods, DINOv2 [84] and Perception Encoder (Percep. Enc.) [85], alongside MLLM-based methods, GPT-4.1 and InternVL3 (8B). As described in Sec. 3.3, Top-1 accuracy (Acc-R@1, %) and mIoU (%) are used for evaluating the spatial and temporal subtasks, respectively. iii) For egocentric depth estimation, we test a general monocular depth model (Depth Anything [79, 80]), a 3D reconstruction-based method (VGGTStream [86, 81]), and two egocentric fisheye-specific models (DAC [87] and UniK3D [88]). For Depth Anything and VGGTStream, input fisheye RGB frames and depth maps are undistorted prior to inference for fair comparison. Additional implementation details (e.g., fps for frame extraction, prompts, and model settings) are provided in the Appendix Sec. A.5.
@@ 4.2 Results on EgoNight-VQA
The main results of all MLLMs are shown in Tab. 1. In addition, we provide per-QA performance comparisons between day (striped bars) and night (solid bars) for paired QA types (Fig. 5(a)) and report nighttime performance across all QA types (Fig. 5(b)), based on averages across all models. Note that non–common case detection is available only in EgoNight-Synthetic, while dynamic events and actions are included only in the real-world data.
From the results in Tab. 1, we observe that almost all MLLMs struggle on our benchmark, with maximum averaged accuracies of 30.93% from the closed-source GPT-4.1, 20.06% from the open-source InternVL3-8B, and 14.29% from the egocentric EgoGPT. Fig. 5(a) further highlights the performance gap, showing declines of 32.8% and 25.0% on EgoNight-Synthetic and EgoNight-Sofia, respectively. Together, these results underscore the substantial challenges posed by our benchmark, exposing the limitations of current MLLMs under nighttime scenarios and highlighting the need for more illumination-robust models. Beyond the overall trends, we note three additional insights from Tab. 1: i) Closed-source models perform best. Within open-source models, Qwen2.5-VL generally improves with scale, yet InternVL outperforms the larger Qwen2.5-VL (72B), suggesting that size alone is insufficient. The relatively low results of EgoGPT further emphasize the need for more robust egocentric models. ii) EgoNight-Oxford achieves the highest scores, but its illumination conditions are more challenging than those in EgoNight-Synthetic and EgoNight-Sofia (Sec. A.4.2, Appendix). This indicates that without paired day videos and our day-augmented auto-labeling strategy, even human annotators face difficulties generating challenging QA pairs, underscoring the practical value of our dataset design; iii) Overall, performance declines across task levels (easy, medium, hard), validating the diversity and difficulty of our benchmark.
Models EgoNight-Synthetic EgoNight-Sofia EgoNight-Oxford Avg.
Easy Medium Hard Easy Medium Hard Easy Medium Hard -
Closed-Source MLLMs
GPT-4.1 29.30 26.87 18.87 32.04 29.35 31.69 39.72 37.13 40.72 30.93
Gemini 2.5 Pro 31.05 24.81 16.51 38.24 26.81 28.87 36.75 36.81 27.88 30.60
Open-source MLLMs
InternVL3-8B 20.21 15.50 16.98 24.03 21.74 20.42 22.90 20.85 16.36 20.06
Qwen2.5-VL-72B 18.39 15.25 12.26 24.03 17.03 20.42 24.81 22.80 16.36 18.99
Qwen2.5-VL-7B 13.01 13.95 13.68 15.44 12.68 12.68 13.74 13.36 12.73 13.44
Qwen2.5-VL-3B 14.69 10.34 7.08 15.50 13.04 12.68 17.18 11.40 12.12 13.41
GLM-4.1V-9B-Base 19.09 13.70 15.57 18.60 18.48 16.20 17.15 22.15 18.79 18.20
VideoLLaMA3-7B 16.85 13.44 14.62 11.11 10.87 9.15 12.26 10.46 9.15 13.64
LLaVA-NeXT-Video-7B 6.36 11.37 1.89 13.95 9.78 14.79 3.05 2.61 3.03 7.28
Egocentric MLLMs
EgoGPT 15.79 13.55 12.04 12.41 12.13 10.36 12.37 13.58 13.68 14.29
Table 1: Comparison results on EgoNight-VQA. Accuracies (%) of OpenQA results across three datasets and three difficulty levels. We compare closed-source models, open-source models, and egocentric-specific models.
@@
Figure 5: Performance analysis of MLLMs on EgoNight-VQA. (a) Day–night performance gap across paired QA types, showing consistent degradation at night. (b) Nighttime performance across all 12 QA types.
@@ 4.3 Results on Day-Night Correspondence Retrieval
The results of day–night retrieval are reported in Tab. 2. The gap between Night–Day and Day–Day shows that cross-illumination retrieval remains highly challenging compared with in-domain retrieval. For spatial retrieval, GPT-4.1 consistently outperforms other methods, achieving over $80\%$ accuracy. This suggests that Retrieval-Augmented Generation methods could further improve performance, as Fig. 5(a) already shows that daytime inputs significantly benefit the models. For temporal retrieval, however, GPT-4.1, despite its strong results on egocentric VQA (Tab. 1) and spatial retrieval, shows a substantial drop compared with feature-based methods (DINOv2 and Perception Encoder). A similar degradation is observed for InternVL3-8B. These findings suggest that while MLLMs excel at spatial semantic understanding, they struggle with temporal reasoning, such as timestamp prediction, which is critical for temporal localization. Further results on temporal limitations are provided in Appendix A.6.
Models Spatial Retrieval (Acc - R@1 $\%$ ↑) Temporal Localization (mIoU $\%$ ↑)
EgoNight-Synthetic EgoNight-Sofia EgoNight-Synthetic EgoNight-Sofia
$toDay Day\toDay Night\toDay$ $toDay Day\toDay Night\toDay$
DINOv2 45.7 28.7 84.5 74.5 - 33.7 - 33.1
Percep. Enc. 65.4 41.6 89.8 80.9 - 32.9 - 33.4
GPT-4.1 75.6 54.1 92.5 84.5 14.7 10.0 21.2 15.5
InternVL3-8B 39.4 27.7 73.9 56.3 10.2 9.9 12.5 13.3
[…]

### v2

@@ 3.2 EgoNight-VQA Benchmark Reconstruction
QA Task Taxonomy. To thoroughly assess models from multiple perspectives, we define a diverse taxonomy of 12 QA tasks. Some of these categories are well-studied and have been explored in previous egocentric VQA benchmarks, such as object/action/text recognition, counting, and spatial reasoning. Others are much less studied or newly proposed in EgoNight-VQA, including scene sequence and navigation (which require not only visual perception but also memory and spatial reasoning), illumination recognition and illumination change (designed to test models’ understanding of lighting concepts), and non–common-sense reasoning (e.g., detecting abnormal cases such as a door inserted into a wall in the synthetic data). More detailed explanations of QA types can be found in the Appendix Sec. LABEL:sec:qa_type_def. We further organize these categories into paired and unpaired QA types, depending on whether the same questions can be consistently applied across day–night counterparts: 1) Paired QA Types. These cover contexts that remain unchanged across day and night, allowing the same QA pairs to be used for both videos and thus providing a clean testbed for measuring performance gaps. Specifically, we include: ① object recognition, ② text recognition, ③ spatial reasoning, ④ scene sequence, ⑤ navigation, ⑥ counting of static, ⑦ action recognition, and ⑧ non–common-sense reasoning. 2) Unpaired QA Types. These include categories that are impractical to pair across day and night, or are only meaningful in the nighttime condition. We consider: ① lighting recognition, ② lighting dynamic, ③ dynamic detection, and ④ counting of dynamic. We control QA clip duration by task type. For static or spatial tasks (e.g., object recognition, lighting recognition), we use short clips of 3 seconds to minimize redundancy; For dynamic or temporal tasks (e.g., action recognition, navigation), the entire video is used to capture the complete context. Following recent works Plizzari et al. (2025); Xiao et al. (2025), we adopt the open-ended QA setting over the closed-form multiple-choice format, as it better reflects natural human–AI interactions.
@@
All three stages are powered by GPT-4.1. Empirically, we find that both the QA-type-specific prompting and the inclusion of daytime videos substantially improve the quality and reliability of the generated QA pairs. More examples of VQA pairs and caption generation can be found in Appendix LABEL:supp:qa_examples and  LABEL:supp:more_visualization.
@@
Dataset Statistics. EgoNight-VQA comprises 3,658 high-quality, fully human-verified QA pairs across 12 task types, sourced from EgoNight-Synthetic, EgoNight-Sofia, and EgoNight-Oxford, with an average of 40 pairs per video. Detailed statistics on QA distribution, video durations, task difficulties, scenarios, and illumination are shown in Fig. 4. The number of videos across the three subsets—Synthetic (50), Sofia (20), and Oxford (20)—is proportionally reflected in the VQA annotations (2029 : 813 : 816). This results in an approximately 1:1 balance between synthetic and real (Sofia + Oxford) VQA samples, ensuring that our benchmark is not dominated by synthetic content. We provide more comparison of Egocentric VQA datasets in Appendix LABEL:sec:dataset_cmp. Overall, EgoNight-VQA provides a diverse and comprehensive benchmark for evaluating egocentric vision models under nighttime conditions.
@@ 3.3 Benchmarks Beyond Egocentric VQA
Day-Night Correspondence Retrieval. To further assess model capabilities beyond VQA, we introduce day–night correspondence retrieval, which evaluates a model’s ability to match visual content across illumination conditions. Specifically, we define two subtasks: i) Spatial Retrieval (Scene Recognition). Spatial retrieval, or scene recognition, is a long-standing vision task Arandjelovic et al. (2016); Miao et al. (2024); Dongfang et al. (2025). Here, it is extended: given a query clip and a set of $N$ candidate clips of equal duration $s$, the model must retrieve the one depicting the same scene. This evaluates a model’s ability to capture and relate spatial relations in egocentric videos, e.g., distinguishing a bedroom from a bathroom or another bedroom. We built this benchmark with 1000 randomly generated meta-tasks. Each task samples a query clip, and the candidate set includes its temporally aligned counterpart (with a temporal shift for added difficulty) plus N1N-1 negatives from other scenes. Performance is measured by Top-1 accuracy across all tasks. In our setup, we use $N=10$, $s=10$ seconds, and a temporal shift of [10, 20] frames. Both Day (query) $\rightarrow$ Day (database) and Day $\rightarrow$ Night settings are evaluated.
@@
ii) Temporal Localization. We further design a temporal localization task to test whether models can align video segments across dynamics. Given a query clip of duration $s$, the model must localize it within the corresponding full video by predicting its start and end timestamps $(t_{i},t_{j})$, directly evaluating temporal reasoning (e.g., grounding “The door is being locked” to 10–20s). We construct 1000 meta-tasks, each generated by randomly sampling one clip from its parent full video that is also randomly selected. Inspired by temporal grounding literature Xin et al. (2024), we adopt mean Intersection-over-Union (mIoU) between the predicted interval $(t_{i},t_{j})$ and the ground-truth interval $(t_{i}^{*},t_{j}^{*})$ as the evaluation metric. Consistent with spatial retrieval, we set $s=10$ seconds and evaluate both Day $\rightarrow$ Day and Night $\rightarrow$ Day settings.
@@
Egocentric Depth Estimation at Night. Depth estimation is a fundamental component of computer vision. On the one hand, extensive research Yang et al. (2024a; b); Wang et al. (2025a) has focused on depth estimation in non-egocentric settings (typically not with fisheye cameras), while egocentric depth estimation remains largely underexplored, especially under nighttime conditions. On the other hand, recent works Chen et al. (2024a); Liu et al. (2025) suggest that incorporating depth can enhance models’ spatial reasoning abilities. These two observations motivate us to construct an auxiliary benchmark for egocentric depth estimation at night. Specifically, we use EgoNight-Synthetic as the testbed, where ground-truth depth maps are provided by the rendering engine. Thanks to the day–night aligned design, we can quantitatively evaluate models under both controlled daytime and nighttime conditions. For evaluation, we adopt standard depth estimation metrics, including absolute relative error (AbsRel), $\delta_{1}(1.25)$, $\delta_{2}(1.25^{2})$, and $\delta_{3}(1.25^{3})$, where $\delta_{k}$ measures the percentage of predicted pixels whose relative error is within a threshold of $1.25^{k}$.
@@ 4.1 Evaluated MLLMs & Metrics
We evaluate a broad set of state-of-the-art MLLMs on the proposed benchmarks. i) For EgoNight-VQA, we include two closed-source commercial models, GPT-4.1 Achiam et al. (2023) and Gemini 2.5 Pro Comanici et al. (2025); eight open-source models, Qwen2.5-VL (3B, 7B, 72B) Bai et al. (2023), VideoLLaMA3 (7B) Zhang et al. (2023b), InternVL3 (8B) Chen et al. (2024b), GLM-4.1V (9B-Base) Hong et al. (2025), and LLaVA-NeXT-Video (7B) Li et al. (2024); as well as EgoGPT Yang et al. (2025), one of the few open-source egocentric models tailored for open-ended QA. Following prior work Plizzari et al. (2025); Fan (2019), we adopt an LLM-as-a-Judge strategy to assess semantic consistency between predictions and ground truth, and report average accuracy across the test sets. ii)We provide further in-depth analysis on synthetic data quality, and potential model improvements. iii) For day–night correspondence retrieval, we benchmark feaFVisture-based retrieval methods, DINOv2 Oquab et al. (2023) and Perception Encoder (Percep. Enc.) Bolya et al. (2025), alongside MLLM-based methods, GPT-4.1 and InternVL3 (8B). As described in Sec. 3.3, Top-1 accuracy (Acc-R@1, %) and mIoU (%) are used for evaluating the spatial and temporal subtasks, respectively. iv) For egocentric depth estimation, we test a general monocular depth model (Depth Anything Yang et al. (2024a; b)), a 3D reconstruction-based method (VGGTStream Zhuo et al. (2025); Wang et al. (2025a)), and two egocentric fisheye-specific models (DAC Guo et al. (2025) and UniK3D Piccinelli et al. (2025)). For Depth Anything and VGGTStream, input fisheye RGB frames and depth maps are undistorted prior to inference for fair comparison. Additional implementation details (e.g., fps for frame extraction, prompts, and model settings) and discussion about LLM-as-a-Judge strategy are provided in the Appendix Sec. LABEL:sec:more_experiment_setup.
@@ 4.2 Results on EgoNight-VQA
The main results of all MLLMs are shown in Tab. LABEL:tab:main_result. In addition, we provide per-QA performance comparisons between day (striped bars) and night (solid bars) for paired QA types (Fig. 5(a)) and report nighttime performance across all QA types (Fig. 5(b)), based on averages across all models. Note that non–common case detection is available only in EgoNight-Synthetic, while dynamic events and actions are included only in the real-world data.
From the results in Tab. LABEL:tab:main_result, we observe that almost all MLLMs struggle on our benchmark, with maximum averaged accuracies of 30.93% from the closed-source GPT-4.1, 20.06% from the open-source InternVL3-8B, and 14.29% from the egocentric EgoGPT. The wide performance spread also confirms that our dataset is sufficiently challenging and effective for distinguishing model capabilities. Fig. 5(a) further highlights the performance gap, showing declines of 32.8% and 25.0% on EgoNight-Synthetic and EgoNight-Sofia, respectively. Together, these results underscore the substantial challenges posed by our benchmark, exposing the limitations of current MLLMs under nighttime scenarios and highlighting the need for more illumination-robust models. Beyond the overall trends, we note three additional insights from Tab. LABEL:tab:main_result: i) Closed-source models perform best. Within open-source models, Qwen2.5-VL generally improves with scale, yet InternVL outperforms the larger Qwen2.5-VL (72B), suggesting that size alone is insufficient. The relatively low results of EgoGPT further emphasize the need for more robust egocentric models. ii) EgoNight-Oxford achieves the highest scores, but its illumination conditions are more challenging than those in EgoNight-Synthetic and EgoNight-Sofia (Sec. LABEL:supp:qa_examples, Appendix). This indicates that without paired day videos and our day-augmented auto-labeling strategy, even human annotators face difficulties generating challenging QA pairs, underscoring the practical value of our dataset design; iii) Overall, performance declines across task levels (easy, medium, hard), validating the diversity and difficulty of our benchmark.
@@
Figure 5: Performance analysis of MLLMs on EgoNight-VQA. (a) Day–night performance gap across paired QA types, showing consistent degradation at night. (b) Nighttime performance across all 12 QA types. NonC means non–common-sense reasoning.
4.3 More in-depth analysis
Quality of EgoNight-Synthetic. More examples of synthetic visualizations are provided in Appendix LABEL:supp:qa_examples and LABEL:sec:more_construction. To quantitatively evaluate how well the synthetic data reflects real-world performance, we benchmark a diverse set of multimodal large language models (MLLMs) on three subsets: Synthetic, Sofia (real), and Oxford (real). For each model, we compute its overall VQA accuracy on each subset and then calculate the Pearson correlation between the average per-model accuracies on synthetic and real-world data (Appendix LABEL:sec:more_exp_result). We observe strong correlations between synthetic and Sofia ($0.9359$, p-value $6.847\times 10^{-5}$), and between synthetic and Oxford ($0.8588$, p-value $1.462\times 10^{-3}$). These results indicate that models that perform well on synthetic data also tend to perform well on real nighttime data, suggesting that the synthetic set preserves the relative difficulty and ranking of models.
To further examine transferability, we fine-tune Qwen2.5-VL-7B using supervised fine-tuning (SFT) on the synthetic training split only, while keeping the real-world data unseen during training. We then evaluate the fine-tuned model directly on the real dataset. The accuracy improves from $14.83\%$ (zero-shot baseline) to $20.57\%$, demonstrating that supervision from synthetic nighttime scenes generalizes to real-world nighttime scenarios.
Impact of Fine-Tuning. To study model adaptation to nighttime egocentric QA, we split EgoNight into $70\%$ training and $30\%$ testing subsets, ensuring no video overlap between splits. We fine-tune Qwen2.5-VL-7B with supervised fine-tuning (SFT) under three configurations: i) full model fine-tuning (updating both vision encoder and LLM), ii) vision-encoder-only tuning (freezing the LLM), and iii) LLM-only tuning (freezing the vision encoder). All models are trained on the same training split and evaluated on the held-out test split.
Tab. LABEL:tab:finetune_all reports the performance for synthetic and real (Sofia and Oxford) separately. It shows consistent performance gains across all configurations compared to the zero-shot baseline. Updating either the vision encoder or the LLM independently improves accuracy, while full fine-tuning achieves the best performance, indicating that both improved visual representations and stronger vision-language alignment contribute to adaptation under low-light conditions.
Perception vs. Reasoning under Fine-Tuning. To better understand which capabilities benefit most from adaptation, we group tasks into perception-oriented (Object Recognition, Text Recognition) and reasoning-oriented (Navigation, Counting) categories and report their accuracies separately in Tab. LABEL:tab:finetune_task.
Perception-oriented tasks show larger absolute improvements after fine-tuning, suggesting that low-level visual adaptation plays a significant role in nighttime scenarios. Vision-encoder-only tuning primarily boosts perception tasks, whereas LLM-only tuning improves both perception and reasoning tasks to a certain extent. However, reasoning-oriented tasks exhibit comparatively smaller gains overall, indicating that high-level reasoning under degraded visual input remains challenging even after adaptation.
Failure Cases. Additional analysis of failure cases is provided in Appendix LABEL:supp:failure_case.
@@ 4.4 Results on Day-Night Correspondence Retrieval
The results of day–night retrieval are reported in Tab. LABEL:tab:retrieval. The gap between Night–Day and Day–Day shows that cross-illumination retrieval remains highly challenging compared with in-domain retrieval. For spatial retrieval, GPT-4.1 consistently outperforms other methods, achieving over $80\%$ accuracy. This suggests that Retrieval-Augmented Generation methods could further improve performance, as Fig. 5(a) already shows that daytime inputs significantly benefit the models. For temporal retrieval, however, GPT-4.1, despite its strong results on egocentric VQA (Tab. LABEL:tab:main_result) and spatial retrieval, shows a substantial drop compared with feature-based methods (DINOv2 and Perception Encoder). A similar degradation is observed for InternVL3-8B. These findings suggest that while MLLMs excel at spatial semantic understanding, they struggle with temporal reasoning, such as timestamp prediction, which is critical for temporal localization. Further results on temporal limitations are provided in Appendix LABEL:sec:more_exp_result.